The following payoff table is given

B3 B4

Find out which option is proposed by the criterion of maximum expected payoff if the probability table of the Bi situations isps % %
Answer #1
=3/15 =2/15 =4/15 =4/15 =2/15
B1 B2 B3 B4 B5 Expected payoff
A1 =B25*B31 =C25*C31 =D25*D31 =E25*E31 =F25*F31 =SUM(B33:F33)
A2 =B26*B31 =C26*C31 =D26*D31 =E26*E31 =F26*F31 =SUM(B34:F34)
A3 =B27*B31 =C27*C31 =D27*D31 =E27*E31 =F27*F31 =SUM(B35:F35)
A4 =B28*B31 =C28*C31 =D28*D31 =E28*E31 =F28*F31 =SUM(B36:F36)
0.2 0.133333 0.26666667 0.266667 0.133333
B1 B2 B3 B4 B5 Expected payoff
A1 0.6 -0.4 2.93333333 2.133333 1.2 6.466667
A2 1.8 0.8 1.6 1.866667 1.6 7.666667
A3 1.6 2.666667 3.2 3.2 1.333333 12
A4 1.8 0.266667 -0.8 1.333333 0.4 3

By the criterion of the maximum payoff value, the option A2 has to be selected and it has the highest payoff value of 7.6667